OpenStudy (anonymous):

What inductance is required with a 17-pF capacitor in order to construct an oscillator capable of generating 550-nm electromagnetic wave. Comment on your answer.

OpenStudy (anonymous):

Okay... do you know anything about LC circuits?

OpenStudy (anonymous):

No, I did not take the course about electricity first sem. so my background in electricity is a bit weak I found out this formula for inductance emf=-Ldi/dt but I am not sure if it is helpful

OpenStudy (anonymous):

Jeez. Okay. Well the physics behind LC circuits is one of the major things you study in an electricity and magnetism course. What class is this question for?

OpenStudy (anonymous):

I am taking an optics course, and it says that electricity and magnetisim is not a prereq.

OpenStudy (anonymous):

i just googled lc circuts, is it just a circut with a capacitor and an inductor?

OpenStudy (anonymous):

Yes it is, but if you haven't studied Electricity I'd be surprised if you were familiar with what either of those things were. Regardless, electricity and magnetism really should be a prerequisite for optics, because light is fundamentally an electromagnetic phenomenon, but oh well. An LC circuit will oscillate with a frequency \[f = \frac{1}{2\pi \sqrt{LC}} \] where L and C are the inductance and capacitance, respectively. You can find the frequency of the electromagnetic wave via the equation \[f\lambda = c\] where lambda and c are the wavelength and speed of light. Just set these things equal to each other and solve for L.

OpenStudy (anonymous):

Wow, thanks a lot man. I can see why I was the only one having trouble with this stuff in class. I will talk to the proff. and see if I need to know other things about electricity for the exams. Amma also be your fan from now on haha, you know your stuff :D cya for now :P

OpenStudy (anonymous):

Thanks, good luck.